Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropriate amount of a 40% antifreeze solution. How many quarts of each solution should she use?
100% antifreeze? qts + 40% solution ? qts = 50% solution qts

Answer by stanbon(57282)

Angela needs 20 quarts of 50% antifreeze solution in her radiator. She plans to obtain this by mixing some pure antifreeze with an appropriate amount of a 40% antifreeze solution. How many quarts of each solution should she use?
100% antifreeze? qts + 40% solution ? qts = 50% solution qts
-----------------------
Let amount of 40% solution be "x"; active ingrediant is 0.40x qts.
Amount of 100% solution is "20-x" qts; active ingrediant is (20-x) qts
Amount of mixed solution is 20 qts ; active ingrediant 0.50*20=10 qts.
EQUATION:
active + active = active
0.4x + 20-x = 10
-0.6x=-10
x= 16 2/3 qts.
She should use 16 2/3 qts. of 40% solution.
Cheers,
Stan H.